OpenStudy (anonymous):

Differentiate. (Assume k is a constant.) y = 1 / (p + ke^p) I tried using the quotient rule... (f/g)' = (gf' - fg') / g^2 and ended up with (1-pke^(p-1)) / (p+ke^p)^2 which is apparently wrong.

OpenStudy (ash2326):

Are we differentiating with respect to p?

OpenStudy (anonymous):

@ash2326 That was all the information I was given :|

OpenStudy (ash2326):

ok, if k is constant then we are differentiating with respect to p. We have e^p in the denominator, that's one place you made a mistake. Do you know differentiation of e^x?

OpenStudy (anonymous):

@ash2326 e^x --> e^x Hmm... standby, going to take another shot at this

OpenStudy (ash2326):

yes, I'm here

OpenStudy (anonymous):

So deriving (p+ke^p) I end up with (1+pke^p) Everything then looks like (0 - (1+pke^p)) / (p+ke^p)^2 Is there any further I can take this?

OpenStudy (anonymous):

Aside from (-1-pke^p) in the numerator

OpenStudy (ash2326):

differentiation of p+ke^p has a small mistake, it should be this \[\frac{d}{dp}(p+ke^P)=1+ke^p\]

OpenStudy (anonymous):

Huh. Looks like I confused myself and mixed the power rule in there. Thanks for all of your help!

OpenStudy (ash2326):

Do you understand?